OpenStudy (anonymous):

Solve the system, using substitution. 3a - b = 11 and 2a + 3b = 0

OpenStudy (anonymous):

I am guessing 3,-2?

OpenStudy (anonymous):

You have to cancel on of the variables first: 3a - b = 11 (multiply by 2 on both sides) 2a +3b = 0 (multiply by 3 on both sides) This makes both of the equations both have 6a. Then you cancel "a" out: 6a-2b = 22 -(6a-9b = 0) ____________ 0a-7b=22 -7b=22 b= -22/7 Subsitute b in for -22/7 in one of the equations, and solve for a. Always check you work!

OpenStudy (anonymous):

That isn't a answer in the multiple choice.

OpenStudy (anonymous):

I made an error: 6a-2b = 22 -(6a-9b = 0) ____________ 0a+7b=22 7b=22 b= 22/7 Is that right?